OpenStudy (anonymous):

There are two regular plots of land ,L1 and L2.The differences between the length and width of L1 is 3m.The length of L1 is 8m more than that of L2 ; The width of L1 is 5m less than that of L2 and the area of L2 is 250m^2 more than half of L1. Find the lengths and widths of L1 and L2.. Help solve please

OpenStudy (anonymous):

area of L1=x^2+3x area of L2=x^2-25 x^2-25=250+1/2(x^2+3X) if you solve the quadratic equation you'll get x=25 dimensions of L1=28 by 25 and L2=30 by 20

OpenStudy (anonymous):

How do you get the area for L1 and L2 ? Can explain?

OpenStudy (anonymous):

Anyone Help =x?

OpenStudy (anonymous):

@juan_chi